# Range, not Independence, Drives Modularity in Biological Inspired Representations

### v2

@@ Appendix C Extension to all Activity Norms
Appendix D Extension to Multidimensional Variables
@@
Appendix E Positive Linear Recurrent Neural Networks
@@
To show this we will first study the weight losses, show that they are minimised if each frequency is modularised from the others. We will then argue that when $\lambda_{W}$ is sufficiently large the representation will contain only two frequencies. Finally, we will study when the activity loss is minimised. Then, as in  Section B.1, when all losses agree on their minima, the representation modularises, else it mixes.
E.1 Structure of Representation - Act I
@@
Where $F$ is smaller than half the number of neurons to ensure the autonomous linear system can propagate each frequency (each frequency ‘uses’ 2 dimensions, as will be obvious from Section E.3.
E.2 Readout Loss
@@
E.3 Recurrent Loss without Bias
@@
${\bm{X}}^{T}{\bm{X}}$ is a symmetric $12F+1\times 2F+$ 1 positive-definite matrix, and its inverse is ${\bm{X}}^{\dagger,T}{\bm{X}}^{\dagger}$, another symmetric positive-definite matrix. To see that these matrices are inverses of one another perform the singular value decomposition, ${\bm{X}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$, and ${\bm{X}}^{\dagger}={\bm{U}}{\bm{S}}^{-1}{\bm{V}}^{T}$. Then ${\bm{X}}^{T}{\bm{X}}={\bm{V}}{\bm{S}}^{2}{\bm{V}}^{T}$ and ${\bm{X}}^{\dagger,T}{\bm{X}}^{\dagger}={\bm{V}}{\bm{S}}^{-2}{\bm{V}}^{T}$, which are clearly inverses of one another.
@@
We then use the following trace inequality, from Ruhe [42]. For two positive semi-definite symmetric matrices, ${\bm{E}}$ and ${\bm{F}}$, with ordered eigenvalues, $e_{1}\geq\ldots\geq e_{n}\geq 0$ and $f_{1}\geq\ldots\geq f_{n}\geq 0$
@@
E.4 Recurrent Loss with Bias
@@
E.5 Structure of Representation - Act II - Large $\lambda_{W}$
@@ E.6 Frequencies
Now consider the case where $\delta>0$ and one of $p$ or $q$ is odd. Then there exists an odd integer $k$ such that[47]:
@@
For any even, odd integer pair of $p,q\neq 1$ and $\delta>0$, $\cos(\pi/p)$ and $\cos(\pi/q)$ are bounded in $(0,1]$. In this case, the maximum possible value of the righthand side is achieved with the smallest possible pair of integers $(3,4)$, for which the inequality does not hold. Thus the inequalities eqn.144, 142 cannot hold at the same time which let us to conclude that there is no $\Delta$ below the critical value satisfying nonnegativity constraint and leads to modular representation.
The last case is where $\delta<0$ and both $p,q$ are odd. We can extend the relation shown in eqn.138 as following by simple substitution of oddity of the variables. With given odd integers $p,q$, there exists an even integer $k$ satisfying eqn.138. Similar to above, take $t=\frac{qk\pi}{p}$ for the mixed encoding eqn.135 and we develop:
@@
, which leads to the following inequality:
@@
Appendix F Metrics for Representational Modularity and Source Statistical Interdependence
@@
where ${\bm{s}}_{-i}$ is a shorthand for $\{{\bm{s}}_{i^{\prime}}\;|\;i^{\prime}\neq i\}$. Since this involves computing mutual information with multiple sources, we restrict ourselves to considering discrete sources and use a continuous-discrete KSG scheme [40] to estimate information with continuous neural activities. We normalise conditional mutual information by $H({\bm{s}}_{i}|{\bm{s}}_{-i})$ to obtain a measure in $[0,1]$. We then arrange the pairwise quantities into a matrix ${\bm{C}}\in\mathbb{R}^{d_{s}\times d_{z}}$ and compute the normalised average “max-over-sum” in a column as a measure of sparsity, following Hsu et al. [22]:
@@
Appendix G What-Where Task
G.1 Experimental Setup
G.1.1 Data Generation
@@ G.1.2 Network Architecture
G.1.3 Training Protocols
The network uses the Adam optimiser ([28]), with learning rate and other hyperparameters varying from experiment to experiment. The mean squared error (MSE) is calculated separately for ‘what’ and ‘where’ tasks, and then combined along with the regularisation terms. The total loss function $L_{total}$ is a combination of the task-specific losses and regularisation terms:
@@ G.1.3 Training Protocols
G.2 Modularity of What & Where
G.2.1 Biological Constraints are necessary for Modularity
@@ G.3 Dropout & Correlation
Appendix H Nonlinear Autoencoders
H.1 Autoencoding Sources
@@ H.2 Autoencoding Images
Appendix I Recurrent Neural Networks
I.1 Linear and Nonlinear Periodic Wave RNN Experiment Details
For linear RNN, we considered a task where the RNN gets a periodic pulse input (2D delta function of frequencies $w_{1},w_{2}$ as input $x$, i.e. $x_{k}(t)=\delta(\cos(w_{k}t)-1),k\in\{1,2\}$) and learns to generate a 2D cosine wave of the same frequencies, $y_{k}(t)=\cos(w_{k}t)$.
For the nonlinear RNN, we designed a two frequency mixing task. Given two source frequencies, the network must generate two cosine waves whose frequencies are the sum and difference of two input frequencies. The network requires non-linearities in order to approximate the multiplication: $\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b),\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)$. In the task, the RNN receives 2D periodic delta pulse of frequency $\frac{w1+w2}{2}$ and $\frac{w1-w2}{2}$ and learns to generate a trajectory of $\cos(w_{1})$ and $\cos(w_{2})$.
@@ I.1 Linear and Nonlinear Periodic Wave RNN Experiment Details
We trained RNN with the trajectory of length $T=200$, bin size $0.1$, and hidden dimension 16. For the linear RNN, we used learning rate 1e-3, 30k training iterations and $\lambda_{\text{target}}=1$, $\lambda_{\text{activity}}=0.5,\lambda_{\text{positivity}}=5,$ and $\lambda_{\text{weight}}=0.02$. For the nonlinear RNN, we used learning rate 7.5e-4, 40k training iterations and $\lambda_{\text{target}}=5,\lambda_{\text{activity}}=0.5$ and $\lambda_{\text{weight}}=0.01$. In both case, we initialised the weights to be orthogonal and the biases at zero, and used Adam optimiser.
To assess the modularity of the trained RNNs, we performed a Fast Fourier Transform(FFT) on each neuron’s activity and measured the relative power of the key frequencies $w1,w2$ with respect to the sum of total power spectrum
@@
and used it as a proxy of mutual information for modularity metric introduced in eqn. 151.
I.2 Nonlinear Teacher-Student RNNs
@@
Appendix J RNN Models of Entorhinal Cortex
We now talk through the linear RNNs used in Section 5. Our model is inspired by the entorhinal literature and linear network models of it [11, 55, 57, 56]. In each trial the agent navigates a 3x3 periodic environment. The RNN receives one special input only at timepoint 0 that tells it the layout of the room (format described later). Otherwise at each timestep it is told which action (north, south, east, west), the agent took. It uses this action to linearly update its hidden state:

### v4

@@ B.4 Extension to all Activity Norms
B.5 Extension to Multidimensional Variables
@@
Appendix C Imperfect Reconstruction
In our results we assume perfect reconstruction. This is not the real situation in either biology or neural networks. If a dataset is range-independent, but only when considering a very unlikely point, then in reality that point might be dropped, incurring a small reconstruction cost, in order to save firing and weight energy.
We illustrate this in the following example. Our dataset comprises four points from the corners of a square - the dataset is range independent. However, we sample these points very nonuniformly, one with probability $\Delta=0.01$, the rest with probability $\Delta=\frac{1-\Delta}{3}$, Figure 7A. If the penalty on reconstruction error is high enough during numerical optimisation, all four points are perfectly reconstructed, this dataset is range independent, and the representation modularises Figure 7B. If the reconstruction loss is made less important the representation changes dramatically, and mixes Figure 7C.
Figure 7: A) We sample a dataset of four points from the corners of a square, one with probability $\Delta$, the rest with probability $\frac{1-\Delta}{3}$. B) If the penalty on reconstruction error is high enough, the data fits perfectly, and the representation is modular. C) If the penalty is lowered, it is preferred to incur reconstruction cost, and mix the representation.
This is an interesting and very relevant effect, that future work could usefully characterise.
Appendix D Positive Linear Recurrent Neural Networks
@@
To show this we will first study the weight losses, show that they are minimised if each frequency is modularised from the others. We will then argue that when $\lambda_{W}$ is sufficiently large the representation will contain only two frequencies. Finally, we will study when the activity loss is minimised. Then, as in Appendix B, when all losses agree on their minima, the representation modularises, else it mixes.
D.1 Structure of Representation - Act I
@@
Where $F$ is smaller than half the number of neurons to ensure the autonomous linear system can propagate each frequency (each frequency ‘uses’ 2 dimensions, as will be obvious from Section D.3.
D.2 Readout Loss
@@
D.3 Recurrent Loss without Bias
@@
${\bm{X}}^{T}{\bm{X}}$ is a symmetric $12F+1\times 2F+$ 1 positive-definite matrix, and its inverse is ${\bm{X}}^{\dagger,T}{\bm{X}}^{\dagger}$, another symmetric positive-definite matrix. To see that these matrices are inverses of one another perform the singular value decomposition, ${\bm{X}}={\bm{U}}\Sigma{\bm{V}}^{T}$, and ${\bm{X}}^{\dagger}={\bm{U}}\Sigma^{-1}{\bm{V}}^{T}$. Then ${\bm{X}}^{T}{\bm{X}}={\bm{V}}\Sigma^{2}{\bm{V}}^{T}$ and ${\bm{X}}^{\dagger,T}{\bm{X}}^{\dagger}={\bm{V}}\Sigma^{-2}{\bm{V}}^{T}$, which are clearly inverses of one another.
@@
We then use the following trace inequality, from Ruhe [Ruhe, 1970]. For two positive semi-definite symmetric matrices, ${\bm{E}}$ and ${\bm{F}}$, with ordered eigenvalues, $e_{1}\geq\ldots\geq e_{n}\geq 0$ and $f_{1}\geq\ldots\geq f_{n}\geq 0$
@@
D.4 Recurrent Loss with Bias
@@
D.5 Structure of Representation - Act II - Large $\lambda_{W}$
@@ D.6 Frequencies
Now consider the case where $\delta>0$ and one of $p$ or $q$ is odd. Then there exists an odd integer $k$ such that Soudry and Speiser [2015]:
@@
For any even, odd integer pair of $p,q\neq 1$ and $\delta>0$, $\cos(\pi/p)$ and $\cos(\pi/q)$ are bounded in $(0,1]$. In this case, the maximum possible value of the righthand side is achieved with the smallest possible pair of integers $(3,4)$, for which the inequality does not hold. Thus the inequalities eqn.96, 94 cannot hold at the same time which let us to conclude that there is no $\Delta$ below the critical value satisfying nonnegativity constraint and leads to modular representation.
The last case is where $\delta<0$ and both $p,q$ are odd. We can extend the relation shown in eqn.90 as following by simple substitution of oddity of the variables. With given odd integers $p,q$, there exists an even integer $k$ satisfying eqn.90. Similar to above, take $t=\frac{qk\pi}{p}$ for the mixed encoding eqn.87 and we develop:
@@
which leads to the following inequality:
@@
Appendix E Metrics for Representational Modularity and Source Statistical Interdependence
@@
where ${\bm{s}}_{-i}$ is a shorthand for $\{{\bm{s}}_{i^{\prime}}\;|\;i^{\prime}\neq i\}$. Since this involves computing mutual information with multiple sources, we restrict ourselves to considering discrete sources and use a continuous-discrete KSG scheme [Ross, 2014] to estimate information with continuous neural activities. We normalise conditional mutual information by $H({\bm{s}}_{i}|{\bm{s}}_{-i})$ to obtain a measure in $[0,1]$. We then arrange the pairwise quantities into a matrix ${\bm{C}}\in\mathbb{R}^{d_{s}\times d_{z}}$ and compute the normalised average “max-over-sum” in a column as a measure of sparsity, following Hsu et al. [2023]:
@@
Appendix F What-Where Task
F.1 Experimental Setup
F.1.1 Data Generation
@@ F.1.2 Network Architecture
F.1.3 Training Protocols
The network uses the Adam optimiser (Kingma and Ba [2014]), with learning rate and other hyperparameters varying from experiment to experiment. The mean squared error (MSE) is calculated separately for ‘what’ and ‘where’ tasks, and then combined along with the regularisation terms. The total loss function $L_{total}$ is a combination of the task-specific losses and regularisation terms:
@@ F.1.3 Training Protocols
F.2 Modularity of What & Where
F.2.1 Biological Constraints are necessary for Modularity
@@ F.3 Dropout & Correlation
Appendix G Nonlinear Autoencoders
G.1 Autoencoding Sources
@@ G.2 Autoencoding Images
Appendix H Recurrent Neural Networks
H.1 Linear and Nonlinear Periodic Wave RNN Experiment Details
For the linear RNN setting we provided a periodic pulse input (2D delta function of frequencies $w_{1},w_{2}$ as input $x$, i.e. $x_{k}(t)=\delta(\cos(w_{k}t)-1),k\in\{1,2\}$) and trained the network to generate a two cosines of the same frequencies, $y_{k}(t)=\cos(w_{k}t)$.
For the nonlinear RNN we designed a two frequency mixing task. Given two source frequencies, the network must generate two cosine waves whose frequencies are the sum and difference of two input frequencies. The network requires non-linearities in order to approximate the multiplication: $\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b),\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)$. In the task we provide the RNN with 2D periodic delta pulse of frequency $\frac{w1+w2}{2}$ and $\frac{w1-w2}{2}$ and learns to generate a trajectory of $\cos(w_{1})$ and $\cos(w_{2})$.
@@ H.1 Linear and Nonlinear Periodic Wave RNN Experiment Details
We trained the RNN with the trajectory of length $T=200$, bin size $0.1$, and hidden dimension 16. For the linear RNN, we used learning rate 1e-3, 30k training iterations and $\lambda_{\text{target}}=1$, $\lambda_{\text{activity}}=0.5,\lambda_{\text{positivity}}=5,$ and $\lambda_{\text{weight}}=0.02$. For the nonlinear RNN, we used learning rate 7.5e-4, 40k training iterations and $\lambda_{\text{target}}=5,\lambda_{\text{activity}}=0.5$ and $\lambda_{\text{weight}}=0.01$. In both case, we initialised the weights to be orthogonal and the biases at zero, and used the Adam optimiser.
To assess the modularity of the trained RNNs, we performed a Fast Fourier Transform(FFT) on each neuron’s activity and measured the relative power of the key frequencies $w_{1},w_{2}$ with respect to the sum of total power spectrum
@@
and used it as a proxy of mutual information for modularity metric introduced in eqn. 103.
H.2 Nonlinear Teacher-Student RNNs
@@
H.3 RNN Models of Entorhinal Cortex
We now talk through the linear RNNs used in Section 5. Our model is inspired by the entorhinal literature and linear network models of it Dorrell et al. [2023], Whittington et al. [2020, 2025, 2023]. In each trial the agent navigates a 3x3 periodic environment. The RNN receives one special input only at timepoint 0 that tells it the layout of the room (format described later). Otherwise at each timestep it is told which action (north, south, east, west), the agent took. It uses this action to linearly update its hidden state via an action-dependent affine transform: